# Chiral Anomaly Induced Transverse Planar Transport Phenomena in Three Dimensional Spin-Orbit Coupled Metals

### v1

@@ II Semiclassical Charge Transport in Boltzmann Transport Equations
Substituting Eqs. (1) and (2) into the kinetic equation (3) and retaining only the term proportional to $(\bm{v}^{s}\!\cdot\!\bm{\Omega}^{s})\bm{B}$ the part responsible for the axial (chiral) anomaly - we obtain
@@ III Chiral Anomaly Induced Charge Transport
where $[dk]\equiv d^{3}k/(2\pi)^{3}$. Employing Eq. (10) we rewrite this as
@@
Hence the current decomposes naturally into its linear and higher‑order parts. The $n$‑th order contribution is
@@ III.1 Linear Transport
Substituting Eqs. (1) and (11) into the $n\!=\!1$ expression of Eq. (15) gives
@@
×$\times\Bigl{[}\frac{e}{\hbar}(\bm{v}^{s}\!\cdot\!\bm{\Omega}^{s})\bm{B}\Bigr{]}\Bigl{(}-\frac{\partial f^{s}_{eq}}{\partial\epsilon^{s}}\Bigr{)}.$
For concreteness we work in the planar geometry $\bm{E}=E_{x}\hat{\bm{x}}$, $\nabla T=\nabla_{x}T\,\hat{\bm{x}}$, and a magnetic field confined to the $x\!-\!y$ plane, $\bm{B}=B\cos\theta\,\hat{\bm{x}}+B\sin\theta\,\hat{\bm{y}}$, with $\theta$ measured from the $x$-axis [Fig.(1)]. We focus on $j^{(1)}_{y}$ as the primary interest of this work lies in the chiral anomaly-induced planar Hall and Nernst effects, both of which manifest through the transverse current. Writing the linear response relations as $\bm{j}^{(1)}_{a}=\sum_{b}\bigl{[}\sigma_{ab}E_{b}-\alpha_{ab}\nabla_{b}T\bigr{]}$ and inserting Eqs. (6) and (7) into Eq. (16) yields
@@ III.2 Nonlinear Transport
×$\hskip 65.00009pt\times\bigl{(}R^{01}_{12}R^{21}_{10}+R^{10}_{10}R^{12}_{12}\bigr{)}\Bigr{]}$
@@ IV.1 The Hamiltonian
[…]
@@ IV.2 Transport Properties
×$\times\left[\frac{(2+\tilde{\mu})\sqrt{1+\tilde{\mu}}}{\tilde{\mu}^{2}}\right]$
@@
×$\times\left[\frac{\tilde{\mu}^{2}+8\tilde{\mu}+8}{\tilde{\mu}^{3}\sqrt{1+\tilde{\mu}}}\right]$
Figure 3: Normalized planar Hall conductivity $\sigma_{yx}/\sigma_{yx}^{0}$ (with $\sigma^{0}_{yx}=\frac{\tau^{*}e^{4}}{8\sqrt{2}\pi^{2}\hbar m^{3/2}\epsilon_{\alpha}^{1/2}}$) as a function of (top) the angle $\theta$ between the magnetic field and current for field strengths $0.5\leq B\leq 1.5$ (color scale), and (bottom) the field magnitude $B$ for angles $40\leq\theta\leq\pi/$ 4 (color scale). We can notice that the maximum value for the field strength occurs when $\theta=\pi/4$.
@@
Although the low‑temperature forms given in Eqs. (42) and (43) satisfy the Mott relation [99] ($\alpha\propto T\partial_{\mu}\sigma$), a closer look at certain fundamental moments—e.g. Eq. (36)—reveals leading‑order terms that decay exponentially in $T$, thereby violating the Mott relation even at relatively low temperatures. Fig. 4 shows the validity of low-temperature expansion for the linear transverse conductivity (Eq.(42)) for varying normalized chemical potential $\mu/\epsilon_{\alpha}$ and temperature $T$. The ratio $\sigma_{yx}/\sigma_{yx}^{\rm LT}$ remains unity (green shading) in the low‑temperature, large‑$\mu/\epsilon_{\alpha}$ regime, indicating agreement with the low temperature expansion. The onset of departure appears as a yellow‑to‑blue gradient, marking where the full calculation deviates from its low‑$T$ approximation. At small $\mu/\epsilon_{\alpha}$, even modest temperatures $T$ suffice to drive the system into the deviation region, whereas at higher $\mu/\epsilon_{\alpha}$ the low‑temperature form holds up to larger $T$. To further understand the validity of the low temperature expansion, Fig. 5 shows a plot comparing the Mott number ($\mathcal{M}=\frac{\alpha}{T\partial_{\mu}\sigma}$) rescaled to unity in the low temperature limit, computed numerically (purple line) with its expected value (green line). A Mott number value close to unity indicates agreement with the Mott relation, while a deviation signals a breakdown of this approximation. We can see that the Mott relation is not valid for higher temperatures and the deviation is apparent at fairly low temperatures (approximately $10\%$ deviation at $20K$ from the ideal value). This violation at relatively low temperatures arises from the exponential behavior of the fundamental moments with respect to temperature (Eqs.(35)-(37)), in contrast to the polynomial dependence assumed in the conventional Mott relation derived via the Sommerfeld expansion [99].
@@
×$\times\left[\frac{12(1+\tilde{\mu})^{\frac{1}{2}}+40(1+\tilde{\mu})^{\frac{3}{2}}+12(1+\tilde{\mu})^{\frac{5}{2}}}{\tilde{\mu}^{5}}\right]$
@@
×$\times\left[\frac{12(1+\tilde{\mu})^{\frac{1}{2}}+40(1+\tilde{\mu})^{\frac{3}{2}}+12(1+\tilde{\mu})^{\frac{5}{2}}}{\tilde{\mu}^{5}}\right]$
@@
×$\times\left[\frac{(1+\tilde{\mu})^{\frac{1}{2}}+(1+\tilde{\mu})^{\frac{3}{2}}}{\tilde{\mu}^{4}}\right]$
Figure 6: Non‑linear normalized planar Hall conductivity $\sigma_{yxx}/\sigma_{yxx}^{0}$ (with $\sigma^{0}_{yxx}=\frac{\tau_{0}\tau^{*}e^{6}}{32\pi^{2}m^{3}\epsilon_{\alpha}^{2}}$) as a function of (top) the in‑plane field angle $\theta$ for $0.5\leq B\leq 1.5$ (color bar), with the linear‑response peak at $\theta=\pi/4$ indicated by a dashed line; and (bottom) the field magnitude $B$ for $40\leq\theta\leq\pi/$ 4 (color bar). Unlike the linear response (Fig. 3), the non‑linear conductivity reaches its maximum at $\theta=\tan^{-1}\!\bigl{(}1/\sqrt{2}\bigr{)}\,<\,\pi/4$
@@ VI Acknowledgment
S.T., R.G.G and B.B.R. acknowledge support from SC Quantum, ARO W911NF2210247 and ONR-N000142312061. G.S. was supported by ANRF-SERB Core Research Grant CRG/2023/005628. G.S. thanks Azaz Ahmad and Gautham Varma K. for useful discussions.
@@ Appendix A CALCULATION OF THE FIRST AND SECOND ORDER PERTURBATION TERMS
$f_{eq}(\epsilon^{s},\mu^{\bar{s}},T^{\bar{s}})=f_{eq}(\epsilon^{s},\mu^{s}-2\delta\mu^{s},T^{s}-2\delta T^{s})$
@@
×$\times\left(-\frac{\partial(f_{eq}^{s}+f^{(1)}_{k}+f^{(2)}_{k}+...)}{\partial\epsilon^{s}}\right)=-\left(\frac{\delta(f^{(1)}_{k}+f^{(2)}_{k}+...)}{\tau^{*}}\right)$
@@
$-\frac{\tau_{\nu}}{2}\left(e\bm{E}.\bm{\Lambda}^{0}_{s}+\frac{\bm{\nabla}T}{T}.\bm{\Lambda}^{1}_{s}\right)=(\delta\mu^{s})D^{0}_{s}+\left(\frac{\delta T^{s}}{T}\right)D^{1}_{s}$ (54)
@@
$-\frac{\tau_{\nu}}{2}\left(e\bm{E}.\bm{\Lambda}^{1}_{s}+\frac{\bm{\nabla}T}{T}.\bm{\Lambda}^{2}_{s}\right)=(\delta\mu^{s})D^{1}_{s}+\left(\frac{\delta T^{s}}{T}\right)D^{2}_{s}$ (55)
@@
×$\times\left(\frac{\partial f_{eq}}{\partial\epsilon^{s}}\right)-2\left(\frac{\tau^{*}}{\tau_{\nu}}\right)\left(\frac{\partial f_{eq}}{\partial\epsilon^{s}}\right)$
×$\times\left[\delta\mu^{s}+\left(\frac{\epsilon^{s}-\mu}{T}\right)\delta T^{s}\right]$ (58)
@@
×$\times\left(-\frac{\partial f^{(1)}_{k}}{\partial\epsilon^{s}}\right)=-\left(\frac{\delta f^{(2)}_{k}}{\tau^{*}}\right)-\frac{2}{\tau_{\nu}}\left[\left(\frac{\partial^{2}f_{eq}}{\partial\epsilon^{s2}}\right)(\delta\mu^{s})^{2}\right.$
$\left.+2\left(\frac{1}{T}\left(\frac{\partial f_{eq}}{\partial\epsilon^{s}}\right)+\left(\frac{\epsilon^{s}-\mu}{T}\right)\left(\frac{\partial^{2}f_{eq}}{\partial\epsilon^{s2}}\right)\right)(\delta\mu^{s})(\delta T^{s})\right]$
$+\frac{2}{\tau_{\nu}}\left[2\left(\frac{\epsilon^{s}-\mu}{T^{2}}\right)\left(\frac{\partial f_{eq}}{\partial\epsilon^{s}}\right)\right.$

### v2

@@ II Semiclassical Charge Transport in Boltzmann Transport Equations
Substituting Eqs. (1) and (2) into the kinetic equation (3) and retaining only the term proportional to $(\bm{v}^{s}\!\cdot\!\bm{\Omega}^{s})\bm{B}$ - the part responsible for the axial (chiral) anomaly - we obtain
@@ III Chiral Anomaly Induced Charge Transport
where $dk]\equiv d^{3}k/(2\pi)^{3}$. Employing Eq. (10) we rewrite this as
@@
Hence the current decomposes naturally into its linear and higher‑order parts. The nn‑th order contribution is
@@ III.1 Linear Transport
Substituting Eqs. (1) and (11) into the n=1n\!=\!1 expression of Eq. (15) gives
@@
$\times\Bigl{[}\frac{e}{\hbar}(\bm{v}^{s}\!\cdot\!\bm{\Omega}^{s})\bm{B}\Bigr{]}\Bigl{(}-\frac{\partial f^{s}_{eq}}{\partial\epsilon^{s}}\Bigr{)}.$
For concreteness we work in the planar geometry $\bm{E}=E_{x}\hat{\bm{x}}$, $\nabla T=\nabla_{x}T\,\hat{\bm{x}}$, and a magnetic field confined to the xyx\!-\!y plane, $\bm{B}=B\cos\theta\,\hat{\bm{x}}+B\sin\theta\,\hat{\bm{y}}$, with $\theta$ measured from the $x$-axis [Fig.(1)]. We focus on $j^{(1)}_{y}$ as the primary interest of this work lies in the chiral anomaly-induced planar Hall and Nernst effects, both of which manifest through the transverse current. Writing the linear response relations as $\bm{j}^{(1)}_{a}=\sum_{b}\bigl{[}\sigma_{ab}E_{b}-\alpha_{ab}\nabla_{b}T\bigr{]}$ and inserting Eqs. (6) and (7) into Eq. (16) yields
@@ III.2 Nonlinear Transport
$\hskip 65.00009pt\times\bigl{(}R^{01}_{12}R^{21}_{10}+R^{10}_{10}R^{12}_{12}\bigr{)}\Bigr{]}$
@@ IV.1 The Hamiltonian
where $s=\pm 1$ labels the two bands, and $k=\mathinner{\!\left\lVert\bm{k}\right\rVert}$. From Eq. (30), we notice that the two bands become degenerate at $\epsilon^{s}=0$. This is also the minimum energy for the $s=+1$ (upper) band. The s=1s=-1 (lower) band reaches its minimum energy at $\epsilon_{\alpha}$, where $\epsilon_{\alpha}=m\alpha^{2}/2\hbar^{2}$. We focus on the regime $\mu\geq 0$, giving each band its own Fermi surface at chemical potential $\mu$. The Chern number for each Fermi surface is
$C_{s}=\iint_{\text{FS}}dS\!\cdot\!\bm{\Omega}^{s}=-s,$ (31)
@@ IV.2 Transport Properties
$\times\left[\frac{(2+\tilde{\mu})\sqrt{1+\tilde{\mu}}}{\tilde{\mu}^{2}}\right]$
@@
$\times\left[\frac{\tilde{\mu}^{2}+8\tilde{\mu}+8}{\tilde{\mu}^{3}\sqrt{1+\tilde{\mu}}}\right$]
Figure 3: Normalized planar Hall conductivity $\sigma_{yx}/\sigma_{yx}^{0}$ (with $\sigma^{0}_{yx}=\frac{\tau^{*}e^{4}}{8\sqrt{2}\pi^{2}\hbar m^{3/2}\epsilon_{\alpha}^{1/2}}$) as a function of (top) the angle $\theta$ between the magnetic field and current for field strengths $0.5\leq B\leq 1.5$ (color scale), and (bottom) the field magnitude $B$ for angles $0\leq\theta\leq\pi/4$ (color scale). We can notice that the maximum value for the field strength occurs when $\theta=\pi/4$.
@@
[…]
@@
$\hskip-8.5359pt\times\left[\frac{12(1+\tilde{\mu})^{\frac{1}{2}}+40(1+\tilde{\mu})^{\frac{3}{2}}+12(1+\tilde{\mu})^{\frac{5}{2}}}{\tilde{\mu}^{5}}\right]$
@@
$\hskip-8.5359pt\times\left[\frac{12(1+\tilde{\mu})^{\frac{1}{2}}+40(1+\tilde{\mu})^{\frac{3}{2}}+12(1+\tilde{\mu})^{\frac{5}{2}}}{\tilde{\mu}^{5}}\right]$
@@
$\times\left[\frac{(1+\tilde{\mu})^{\frac{1}{2}}+(1+\tilde{\mu})^{\frac{3}{2}}}{\tilde{\mu}^{4}}\right]$
Figure 6: Non‑linear normalized planar Hall conductivity $\sigma_{yxx}/\sigma_{yxx}^{0}$ (with $\sigma^{0}_{yxx}=\frac{\tau_{0}\tau^{*}e^{6}}{32\pi^{2}m^{3}\epsilon_{\alpha}^{2}}$) as a function of (top) the in‑plane field angle $\theta$ for $0.5\leq B\leq 1.5$ (color bar), with the linear‑response peak at $\theta=\pi/4$ indicated by a dashed line; and (bottom) the field magnitude $B$ for $0\leq\theta\leq\pi/4$ (color bar). Unlike the linear response (Fig. 3), the non‑linear conductivity reaches its maximum at $\theta=\tan^{-1}\!\bigl{(}1/\sqrt{2}\bigr{)}\,<\,\pi/4$
@@ VI Acknowledgment
S.T., R.G.G and B.B.R. acknowledge support from SC Quantum, ARO W911NF2210247 and ONR-N000142312061. BBR acknowledges support from the National Center for Transportation Cybersecurity and Resiliency (TraCR) (a US Department of Transportation National University Transportation Center) headquartered at Clemson University, Clemson, South Carolina, USA. Any opinions, findings, conclusions, and recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of TraCR. The US Government assumes no liability for the contents or use thereof. G.S. was supported by ANRF-SERB Core Research Grant CRG/2023/005628. G.S. thanks Azaz Ahmad and Gautham Varma K. for useful discussions.
@@ Appendix A CALCULATION OF THE FIRST AND SECOND ORDER PERTURBATION TERMS
$\hskip-28.45274ptf_{eq}(\epsilon^{s},\mu^{\bar{s}},T^{\bar{s}})=f_{eq}(\epsilon^{s},\mu^{s}-2\delta\mu^{s},T^{s}-2\delta T^{s})$
@@
$\times\left(-\frac{\partial(f_{eq}^{s}+f^{(1)}_{k}+f^{(2)}_{k}+...)}{\partial\epsilon^{s}}\right)=-\left(\frac{\delta(f^{(1)}_{k}+f^{(2)}_{k}+...)}{\tau^{*}}\right)$
@@
$\hskip-28.45274pt-\frac{\tau_{\nu}}{2}\left(e\bm{E}.\bm{\Lambda}^{0}_{s}+\frac{\bm{\nabla}T}{T}.\bm{\Lambda}^{1}_{s}\right)=(\delta\mu^{s})D^{0}_{s}+\left(\frac{\delta T^{s}}{T}\right)D^{1}_{s}$ (54)
@@
$\hskip-28.45274pt-\frac{\tau_{\nu}}{2}\left(e\bm{E}.\bm{\Lambda}^{1}_{s}+\frac{\bm{\nabla}T}{T}.\bm{\Lambda}^{2}_{s}\right)=(\delta\mu^{s})D^{1}_{s}+\left(\frac{\delta T^{s}}{T}\right)D^{2}_{s}$ (55)
@@
$\times\left(\frac{\partial f_{eq}}{\partial\epsilon^{s}}\right)-2\left(\frac{\tau^{*}}{\tau_{\nu}}\right)\left(\frac{\partial f_{eq}}{\partial\epsilon^{s}}\right)$
$\times\left[\delta\mu^{s}+\left(\frac{\epsilon^{s}-\mu}{T}\right)\delta T^{s}\right]$ (58)
@@
$D\left(\bm{v}^{s}+\frac{e}{\hbar}(\bm{v}^{s}\cdot\bm{\Omega}^{s})\bm{B}\right)\cdot\left(e\bm{E}+\left(\frac{\epsilon^{s}-\mu}{T}\right)\bm{\nabla}T\right)\ \left(-\frac{\partial f^{(1)}_{k}}{\partial\epsilon^{s}}\right)=-\left(\frac{\delta f^{(2)}_{k}}{\tau^{*}}\right)-\frac{2}{\tau_{\nu}}\left[\left(\frac{\partial^{2}f_{eq}}{\partial\epsilon^{s2}}\right)(\delta\mu^{s})^{2}\right$.
$\left.+2\left(\frac{1}{T}\left(\frac{\partial f_{eq}}{\partial\epsilon^{s}}\right)+\left(\frac{\epsilon^{s}-\mu}{T}\right)\left(\frac{\partial^{2}f_{eq}}{\partial\epsilon^{s2}}\right)\right)(\delta\mu^{s})(\delta T^{s})\right]+\frac{2}{\tau_{\nu}}\left[2\left(\frac{\epsilon^{s}-\mu}{T^{2}}\right)\left(\frac{\partial f_{eq}}{\partial\epsilon^{s}}\right)\right.\left.+\left(\frac{\epsilon^{s}-\mu}{T}\right)^{2}\left(\frac{\partial^{2}f_{eq}}{\partial\epsilon^{s2}}\right)\right](\delta T^{s})^{2}$ (59)